I'm having trouble helping my daughter solve two step equations on the following two : f-6 over the number 5 =3.2 then
44= -4 + 8p

Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

f-6 over the number 5 =3.2
-------------
(f-6)/5 = 3.2
f-6=5(3.2)
f-6=16
f=22
-------------
then
44= -4 + 8p
48 = 8p
p=6
---------
Cheers,
Stan H.

Answer by checkley75(3666)

(F-6)/5=3.2 CROSS MULTIPLY THUS
F-6=5*3.2
F-6=16
F=16+6
F=22 ANSWER
PROOF
(22-6)/5=3.2
16/5=3.2
3.2=3.2